1. A pl.astic bO% 1.6 m long, 1.25 m wUk and 65 cm deep is to be made. It is opened at the top. Ignori ng the thickness of the plastic sheet, determine:
(i) The area of the slz:eet required for making the box. ;

182
(ii) The cost of sheet for it, if a sheet measuring 1 m2 costs f 20.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
. Here, l = 1.6 m, b = 1.26 m, h = 65 cm = 0.65 m. (i) Area = 2(1 + b) x h + lb = (2(1.6 + 1.25) x 0.66 + 1.6 x 1.25) m2 = (2 x 2.76 x 0.65 + 1.5 x 1.25) m2 = (3.575 + 1.875) m2 = 5.45 m2 (ii) Cost = l' 20 x 5.45 = l' 109.